Consider the titration of 43.4 mL of 0.265 M HF with 0.195 M NaOH. Calculate the pH at each of the following points.

a. How many milliliters of base are required to reach the equivalence point?

b Calculate the pH after the addition of 10.9 mL of base

c. Calculate the pH at halfway to the equivalence point.

d. Calculate the pH at the equivalence point.

e. Calculate the pH after the addition of 86.8 mL of base.
Answer #1

Given that Molarity of HF 0.265 M Volume of HF = 43.4 mL Ka of HF= 3.5 x 104 Concentration of NaOH 0.195 M (a) Number of mole

b) Addition of 10.9 mL of NaOH Number of moles of HF= M*V = 0.265 M* 43.4 mL 11.501 mmol Number of moles of NaOH = M*V = 0.19

c) At half equivalence point, the amount of base formed is equal to the amount of acid unreacted So, pH pKa 3.46

d) At equivalence point: Volume of NaOH required to reach equivalence point 58.98 mL This is equivalence point. At this point

х.х К, 0.1123-x -2.86x10-11 x is small, so 0.1123-x= 0.1123 x23.21x10-12 x=1.79x10 [OH 1.79x10M pOH log[OH ]=-log (1.79x10) =